# Physics-based Full-band GaN HEMT Simulation Suggests Upper Bound of LO Phonon Lifetime

### v1

@@ Abstract
Intrinsic limits to device performance arise from fundamental material properties that define the best achievable operation, independent of engineering constraints. In GaN HEMTs, hot longitudinal optical (LO) phonons can act as an intrinsic performance bottleneck by reducing electron saturation velocity, output current, and transconductance–metrics that are important for device operation. While bulk GaN studies report LO phonon lifetimes of $\sim$1 ps, leading to strong nonequilibrium phonon populations, ungated heterostructures show much shorter lifetimes of only tens of femtoseconds. Because direct measurement in HEMTs is challenging, the true impact of hot phonons remains uncertain. Full-band transport simulations of a fabricated GaN HEMT presented here reveal that LO phonon lifetimes must be $\lesssim$ 40 fs to reproduce measured I–V characteristics, consistent with ultrafast decay observed in GaN heterostructures. We show that even this ultrafast LO-phonon decay is insufficient to fully suppress hot-phonon effects: the residual nonequilibrium LO population continues to limit the current density at high bias. Moreover, when the LO-phonon lifetime exceeds a few tens of femtoseconds, a pronounced hot-phonon bottleneck emerges, leading to a substantial current-density suppression that is inconsistent with experimental observations.
@@ I Introduction
GaN-based high-electron-mobility transistor (HEMT) is an attractive technology for high-frequency and high-power electronics needed in satellite communications, radar systems, and wireless communication networks. hoo2021emerging; lu2025review The exceptional material properties of GaN, including its wide bandgap, high breakdown voltage, meneghesso2014breakdown and large peak electron velocity, barker2005bulk enable devices to operate under high electric fields and deliver superior power densities. The strong spontaneous and piezoelectric polarizations in AlGaN/GaN heterostructures give rise to a high-density two-dimensional electron gas (2DEG) at the interface without the need for modulation doping, ambacher1999two an intrinsic feature that forms the basis for the excellent electrical behavior observed in GaN HEMTs.
@@
Electron-phonon scatterings in a device often set intrinsic limits on device performance, independent of engineering or processing constraints. In the case of GaN, the optical phonon energy is high because of the light mass of the nitrogen atom, which combined with the high electronegativity of nitrogen, results in a high electron-LO phonon scattering rate. khurgin2007hot If a hot electron in GaN relaxes its energy by emitting an LO phonon that does not decay rapidly into acoustic modes, it could lead to the accumulation of hot LO phonons srivastava2008origin; barman2004long; Ridley_1996; ridley2004; ridley2013quantum since the velocity of optical phonons is negligible. matulionis2006hot; fang2012effect This accumulation of hot LO phonons could result in an intrinsic performance bottleneck of the GaN technology, necessitating innovations in heat extraction and phonon engineering. This paper’s central goal is to clearly answer the question of whether hot LO phonons are critical to the performance of a GaN HEMT and compare our predictions of hot LO phonon lifetimes to prior measurements in ungated GaN heterostructures.
@@
Electron transport in a fabricated GaN HEMT is simulated using our in-house deterministic full-band Boltzmann solver, Fermi Kinetics Transport (FKT), 10.1063/1.3270404; grupen2011energy; 7505603; tunga2022comparison which explicitly captures LO phonon dynamics. Simulations compared to measured I–V characteristics confirm that, unlike bulk GaN where LO phonon lifetimes can be relatively long, the effective lifetime in GaN HEMTs is significantly shorter. This suppression of phonon lifetime is attributed to plasmon–phonon coupling and the influence of heterointerfaces inherent to HEMT structures. Ridley_Dyson_2009; dyson2011lifetime Additional simulations show that, even with ultrafast LO-phonon decay, the hot-phonon bottleneck in GaN HEMTs persists, leading to as much as 30-60% reduction in both saturation current and peak transconductance, reinforcing that LO phonons remain an intrinsic performance limiter. Moreover, LO-phonon lifetimes exceeding a few tens of femtoseconds produce phonon heating and I–V characteristics that deviate markedly from experimental measurements, indicating that such long lifetimes are not physically realistic for GaN HEMTs.
@@ II Fermi Kinetics Transport Solver
Our work is based on FKT, which solves the Boltzmann transport equation (BTE) in the deterministic sense, making it computationally more efficient than Monte Carlo solvers. Unlike other deterministic Boltzmann solvers, FKT uniquely handles electronic heat flow using the heat capacity of an ideal Fermi gas instead of relying on an arbitrarily defined electron thermal conductivity. This alternative closure relation enhances the numerical stability and convergence of FKT compared to traditional solvers. 10.1063/1.3270404 FKT employs the method of moments to compute both electron and energy fluxes and is capable of incorporating detailed electronic bandstructures and key scattering mechanisms based on Fermi’s golden rule. grupen2011energy It can also couple rapid hot electron dynamics with much slower defect ionization processes to simulate operational instabilities and with full wave electromagnetic fields to enable large signal RF simulations. 7548427 These features make FKT a versatile tool for advanced electron transport modeling.
@@ II.1 Incorporating Electronic Band Structure and Scattering
where $f_{i}(E)$ is the Fermi-Dirac distribution function assigned to the $i$-th local carrier population, and the term in square brackets is the density of electron states $\rho_{\mathbf{k}}(E)$ integrated over all degenerate momentum vectors associated with the local distribution function $f_{i}(E)$. Since multiple Fermi-Dirac distributions are assigned to a given point in real space, the total density is obtained by summing the contributions. From the first moment of the BTE, the particle flux is given by an integral of the form:
@@
$J_{i}=\int_{E_{i}}\underbrace{\left[\int_{\mathbf{k}_{i}}\mathbf{v}\,\mathbf{v}^{T}\,\tau_{\mathbf{k}}\,\rho_{\mathbf{k}}(E)\,d\mathbf{k}\right]}_{\mu}\left(qE\,\frac{\partial f_{i}}{\partial E}-\nabla f_{i}\right)dE,$ (2)
where electron velocity $\mathbf{v}=\nabla_{\mathbf{k}}E/\hbar$, and $\tau_{\mathbf{k}}(E)$ is the momentum relaxation time contributed by all scattering mechanisms, while $\mu$ represents the energy isosurface integral. 7505603
@@
The overview of the entire preprocessing stage is illustrated in Fig. 1. An empirical pseudopotential method (EPM)-computed band structure, 6022750 shown in Fig. 1(a), is employed for wurtzite GaN, and the irreducible wedge of its first Brillouin zone, shown in Fig. 1(b), is filled with a conformal tetrahedral mesh to determine the electron energies at discrete $\mathbf{k}$-points. Using piecewise-linear interpolation along the mesh edges, the electron energy isosurfaces are extracted grupen2011energy as shown in Fig. 1(c). The integral $\mu$ in Eq. (2) is computed over each energy isosurface to produce an energy spectrum like that shown in Fig. 1(d). Because lattice temperature, LO phonon temperature, ionized defect density, and electron density vary sensitively during device operation, the flux isosurface integrals must evolve correspondingly. In order to capture the dependence of these four quantities, the computed integrals are first fitted to polynomials of electron energy, and the resulting polynomial coefficients are subsequently parameterized as functions of the four varying quantities, i.e., lattice temperature, LO phonon temperature, ionized defect density, and electron density. Therefore, the simulator is capable of using these four changing quantities to obtain the correct flux isosurface integrals matching the actual numerical integrals.
@@
Figure 2 highlights the necessity to explicitly consider LO phonon temperature in the solver. This figure plots the $\Gamma$ valley isosurface integrals when the LO phonon temperature is varied while the other parameters (lattice temperature, ionized defect density, and electron density) are held constant. The energy isosurface integral increases with energy up to approximately 92 meV, driven by the growth of the velocity vector $\mathbf{v}$ and the density of states $\rho_{\mathbf{k}}(E)$, even though $\tau_{\mathbf{k}}(E)$ itself begins to decline. At around 92 meV, when electrons begin to emit optical phonons in addition to absorbing them, $\tau_{\mathbf{k}}(E)$ decreases sharply, leading to a reduction in the overall integral. At higher energies, the integrals gradually rise again. With increasing LO phonon temperature, the phonon occupation number rises, enhancing both absorption and emission scattering rates. Consequently, $\tau_{\mathbf{k}}(E)$ decreases across all electron energies, leading to a uniform reduction in the flux isosurface integrals throughout the spectrum.
@@ II.2 Treatment of Acoustic and Optical Phonons
where $U_{\mathrm{LO}}$ is the LO phonon collision operator $\hbar\omega_{\rm LO}U_{\rm LO}=C^{\mathrm{E}}$, $n_{\mathrm{q}}$ is the phonon number, and $n_{\mathrm{q}}^{0}$ is the LO phonon occupation number when its temperature $T_{\rm LO}=T_{\rm A}$.
For simplicity, $n_{\mathrm{q}}$ in Eq. (4) is assumed symmetric in $\mathbf{q}$-space. To estimate $|\mathbf{q}|_{\mathrm{max}}$ for hot LO phonons, a series of bulk GaN electron drift velocity versus electric field simulations, like those shown in Fig. 3, were conducted. As indicated in Fig. 4, these simulations also computed LO phonon occupation numbers for different ranges of $|\mathbf{q}|$. Based on these data, $|\mathbf{q}|_{\mathrm{max}}=6\times 10^{6}$ 1/cm was used in Eq. (4) for GaN HEMT device simulations.
@@ III.2 Device Simulation
Measured and simulated GaN HEMT drain currents $I_{\mathrm{D}}$ versus drain voltage $V_{\mathrm{DS}}$ for different gate biases $V_{\mathrm{GS}}$ are shown in Fig. 6. The favorable comparison between measured and simulated data indicates that FKT can accurately capture the physics of the fabricated device. Marino2010 The solver predicts that the LO phonon lifetime must be as short as few tens of fs in order to explain the measured I–V characteristics. Specifically, for the results in Fig. 6, $\tau_{\mathrm{LO}}=30$ fs. While such a short phonon lifetime may seem unexpected for bulk GaN, it is reasonable in GaN heterostructures, where LO phonon interactions with interfaces create additional decay channels that enable rapid conversion into heat-conducting modes. matulionis2009ultrafast; liberis2014hot The impact of $\tau_{\mathrm{LO}}$ on the output current at $V_{\mathrm{GS}}=1$ V and $V_{\mathrm{DS}}=10$ V is shown in Fig. 7. As expected, increasing $\tau_{\mathrm{LO}}$ beyond 30 fs reduces the drain current. Conversely, decreasing $\tau_{\mathrm{LO}}$ from 30 fs to 1 fs yields a $\sim$30% increase in drain current, indicating that even at ultrashort $\tau_{\mathrm{LO}}$ the device performance remains constrained by the hot-LO-phonon bottleneck.
Figure 6: Simulated output curves with 30 fs hot LO phonon lifetime. Symbols are experimental data from Marino et al. Marino2010
@@
The role of LO phonon heating is further illustrated in Fig. 8(a), which presents the simulated output characteristics at $V_{\mathrm{GS}}=1$ V under four different conditions. The simulation with $\tau_{\mathrm{LO}}=30$ fs is the baseline result (labeled as “All”), which matches well with the measured data shown in Fig. 6. When LO phonon heating with $\tau_{\mathrm{LO}}=30$ fs is considered but acoustic phonon heating is ignored (result labeled as “LO”), the drain current merely increases by 5% at $V_{\mathrm{GS}}=1$ V and $V_{\mathrm{DS}}=10$ V, compared to the baseline case. This observation reinforces that LO phonon heating alone introduces substantial performance degradation even in the absence of acoustic-phonon-induced self-heating. When we assume an instantaneous decay of LO phonons into acoustic modes, effectively removing the hot-LO phonon contribution and considering only longitudinal and transverse acoustic phonon heating (labeled as “LA/TA”), $I_{\mathrm{D}}$ increases by approximately 30% at $V_{\mathrm{GS}}=1$ V and $V_{\mathrm{DS}}=10$ V. In contrast, eliminating both hot LO and hot acoustic phonons in the device (labeled as “None”) would increase the device current by up to 50%, suggesting that phonon-engineering strategies could, in principle, enhance intrinsic device performance. However, we are not aware of any studies demonstrating LO-phonon lifetimes below $30\pm 15$ fs in GaN heterostructures, indicating that such aggressive lifetime reduction has not yet been realized experimentally.

### v3

# Physics-based Full-band GaN High-Electron-Mobility Transistor Simulation Suggests Upper Bound of LO Phonon Lifetime

@@ Abstract
Intrinsic limits to device performance arise from fundamental material properties that define the best achievable operation, independent of engineering constraints. In GaN high-electron-mobility transistors (HEMTs), hot longitudinal optical (LO) phonons can act as an intrinsic performance bottleneck by reducing electron saturation velocity, output current, and transconductance–metrics that are important for device operation. While bulk GaN studies report LO phonon lifetimes of $\sim$1 ps, leading to strong nonequilibrium phonon populations, ungated heterostructures show much shorter lifetimes of only tens of femtoseconds. Because direct measurement in HEMTs is challenging, the true impact of hot phonons remains uncertain. Full-band transport simulations of a fabricated GaN HEMT presented here reveal that LO phonon lifetimes must be $\lesssim$ 40 fs to reproduce measured I–V characteristics, consistent with ultrafast decay observed in GaN heterostructures. We show that even this ultrafast LO-phonon decay is insufficient to fully suppress hot-phonon effects: the residual nonequilibrium LO population continues to limit the current density at high bias. Moreover, when the LO-phonon lifetime exceeds a few tens of femtoseconds, a pronounced hot-phonon bottleneck emerges, leading to a substantial current-density suppression that is inconsistent with experimental observations.
@@ I Introduction
GaN-based high-electron-mobility transistor (HEMT) is an attractive technology for high-frequency and high-power electronics needed in satellite communications, radar systems, and wireless communication networks. Hoo Teo et al. (2021); Lu et al. (2025) The exceptional material properties of GaN, including its wide bandgap, high breakdown voltage, Meneghesso et al. (2014) and large peak electron velocity, Barker et al. (2005) enable devices to operate under high electric fields and deliver superior power densities. The strong spontaneous and piezoelectric polarizations in AlGaN/GaN heterostructures give rise to a high-density two-dimensional electron gas (2DEG) at the interface without the need for modulation doping, Ambacher et al. (1999) an intrinsic feature that forms the basis for the excellent electrical behavior observed in GaN HEMTs.
@@
Electron-phonon scatterings in a device often set intrinsic limits on device performance, independent of engineering or processing constraints. In the case of GaN, the optical phonon energy is high because of the light mass of the nitrogen atom, which combined with the high electronegativity of nitrogen, results in a high electron-LO phonon scattering rate. Khurgin et al. (2007) If a hot electron in GaN relaxes its energy by emitting an LO phonon that does not decay rapidly into acoustic modes, it could lead to the accumulation of hot LO phonons Srivastava (2008); Barman and Srivastava (2004); Ridley (1996); Ridley et al. (2004); Ridley FRS (2013) since the velocity of optical phonons is negligible. Matulionis (2006); Fang et al. (2012) This accumulation of hot LO phonons could result in an intrinsic performance bottleneck of the GaN technology, necessitating innovations in heat extraction and phonon engineering. This paper’s central goal is to clearly answer the question of whether hot LO phonons are critical to the performance of a GaN HEMT and compare our predictions of hot LO phonon lifetimes to prior measurements in ungated GaN heterostructures.
@@
Electron transport in a fabricated GaN HEMT is simulated using our in-house deterministic full-band Boltzmann solver, Fermi Kinetics Transport (FKT), Grupen (2009, 2011, 2016); Tunga et al. (2022) which explicitly captures LO phonon dynamics. Simulations compared to measured I–V characteristics confirm that, unlike bulk GaN where LO phonon lifetimes can be relatively long, the effective lifetime in GaN HEMTs is significantly shorter. This suppression of phonon lifetime is attributed to plasmon–phonon coupling and the influence of heterointerfaces inherent to HEMT structures. Ridley and Dyson (2009); Dyson and Ridley (2011) Additional simulations show that, even with ultrafast LO-phonon decay, the hot-phonon bottleneck in GaN HEMTs persists, leading to as much as 30-60% reduction in both saturation current and peak transconductance, reinforcing that LO phonons remain an intrinsic performance limiter. Moreover, LO-phonon lifetimes exceeding a few tens of femtoseconds produce phonon heating and I–V characteristics that deviate markedly from experimental measurements, indicating that such long lifetimes are not physically realistic for GaN HEMTs.
@@ II Fermi Kinetics Transport Solver
Our work is based on FKT, which solves the Boltzmann transport equation (BTE) in the deterministic sense, making it computationally more efficient than Monte Carlo solvers. Unlike other deterministic Boltzmann solvers, FKT uniquely handles electronic heat flow using the heat capacity of an ideal Fermi gas instead of relying on an arbitrarily defined electron thermal conductivity. This alternative closure relation enhances the numerical stability and convergence of FKT compared to traditional solvers. Grupen (2009) FKT employs the method of moments to compute both electron and energy fluxes and is capable of incorporating detailed electronic bandstructures and key scattering mechanisms based on Fermi’s golden rule. Grupen (2011) It can also couple rapid hot electron dynamics with much slower defect ionization processes to simulate operational instabilities and with full wave electromagnetic fields to enable large signal RF simulations. Miller et al. (2016) These features make FKT a versatile tool for advanced electron transport modeling.
@@ II.1 Incorporating Electronic Band Structure and Scattering
where $f_{\mathrm{i}}(E)$ is the Fermi-Dirac distribution function assigned to the $i$-th local carrier population, and the term in square brackets is the density of electron states $\rho_{\mathrm{k}}(E)$ integrated over all degenerate momentum vectors associated with the local distribution function $f_{\mathrm{i}}(E)$. Since multiple Fermi-Dirac distributions are assigned to a given point in real space, the total density is obtained by summing the contributions. From the first moment of the BTE, the particle flux is given by an integral of the form:
@@
$J_{\mathrm{i}}=\int_{E_{\mathrm{i}}}\underbrace{\left[\int_{\mathbf{k}_{\mathrm{i}}}\mathbf{v}\,\mathbf{v}^{T}\,\tau_{\mathrm{k}}\,\rho_{\mathrm{k}}(E)\,d\mathbf{k}\right]}_{\mu}\left(qE\,\frac{\partial f_{\mathrm{i}}}{\partial E}-\nabla f_{\mathrm{i}}\right)dE,$ (2)
where electron velocity $\mathbf{v}=\nabla_{\mathbf{k}}E/\hbar$, and $\tau_{\mathrm{k}}(E)$ is the momentum relaxation time contributed by all scattering mechanisms, while $\mu$ represents the energy isosurface integral. Grupen (2016) Note that electrostatic screening is included in the charged impurity and piezoelectric scattering mechanisms but is neglected in the optical phonon scattering processes. The collision operator for polar optical phonon scattering in our simulations is derived from the Fröhlich Hamiltonian. Although screening is not explicitly included in this operator at present, we believe this to be a reasonable approximation because preliminary estimates of the Thomas-Fermi length Go et al. (2025) indicate that screening may reduce our collision operator by, at most, a factor of 4 and would likely be even less significant in the HEMT’s access regions where the collision operator has its greatest impact on device behavior. Nevertheless, as part of ongoing improvements to the solver, we plan to further investigate screening of polar optical phonon collision operator.
@@
The overview of the entire preprocessing stage is illustrated in Fig. 1. An empirical pseudopotential method (EPM)-computed band structure, Sridharan et al. (2011) shown in Fig. 1(a), was employed for wurtzite GaN, and the irreducible wedge of its first Brillouin zone, shown in Fig. 1(b), was filled with a conformal tetrahedral mesh to determine the electron energies at discrete $\mathbf{k}$-points. Using piecewise-linear interpolation along the mesh edges, the electron energy isosurfaces are extracted Grupen (2011) as shown in Fig. 1(c). The integral $\mu$ in Eq. (2) is computed over each energy isosurface to produce an energy spectrum like that shown in Fig. 1(d). Because lattice temperature, LO phonon temperature, ionized defect density, and electron density vary sensitively during device operation, the flux isosurface integrals must evolve correspondingly. In order to capture the dependence of these four quantities, the computed integrals were first fitted to polynomials of electron energy, and the resulting polynomial coefficients were subsequently parameterized as functions of the four varying quantities, i.e., lattice temperature, LO phonon temperature, ionized defect density, and electron density. Therefore, the simulator is capable of using these four changing quantities to obtain the correct flux isosurface integrals matching the actual numerical integrals.
@@
Figure 2 highlights the necessity to explicitly consider LO phonon temperature in the solver. This figure plots the $\Gamma$ valley isosurface integrals when the LO phonon temperature is varied while the other parameters (lattice temperature, ionized defect density, and electron density) are held constant. The energy isosurface integral increases with energy up to approximately 92 meV, driven by the growth of the velocity vector $\mathbf{v}$ and the density of states $\rho_{\mathbf{k}}(E)$, even though $\tau_{\mathrm{k}}(E)$ itself begins to decline. At around 92 meV, when electrons begin to emit optical phonons in addition to absorbing them, $\tau_{\mathrm{k}}(E)$ decreases sharply, leading to a reduction in the overall integral. At higher energies, the integrals gradually rise again. With increasing LO phonon temperature, the phonon occupation number rises, enhancing both absorption and emission scattering rates. Consequently, $\tau_{\mathrm{k}}(E)$ decreases across all electron energies, leading to a uniform reduction in the flux isosurface integrals throughout the spectrum.
@@ II.2 Treatment of Acoustic and Optical Phonons
where $U_{\mathrm{LO}}$ is the LO phonon collision operator $\hbar\omega_{\rm LO}U_{\rm LO}=C^{\mathrm{E}}$, $n_{\mathrm{q}}$ is the phonon number, and $n_{\mathrm{q}}^{0}$ is the LO phonon occupation number when its temperature $T_{\rm LO}=T_{\rm A}$. For simplicity, $n_{\mathrm{q}}$ in Eq. (4) is assumed symmetric in $\mathbf{q}$-space. To estimate $|\mathbf{q}|_{\mathrm{max}}$ for hot LO phonons, a series of bulk GaN electron drift velocity versus electric field simulations, like those shown in Fig. 3, were conducted. As indicated in Fig. 4, these simulations also computed LO phonon occupation numbers for different ranges of $|\mathbf{q}|$. Based on these data, $|\mathbf{q}|_{\mathrm{max}}=6\times 10^{6}$ 1/cm was used in Eq. (4) for GaN HEMT device simulations.
@@ III.2 Device Simulation
Measured and simulated GaN HEMT drain currents $I_{\mathrm{D}}$ versus drain voltage $V_{\mathrm{DS}}$ for different gate biases $V_{\mathrm{GS}}$ are shown in Fig. 6. The favorable comparison between measured and simulated data indicates that FKT can accurately capture the physics of the fabricated device. Marino et al. (2010) The solver predicts that the LO phonon lifetime must be as short as few tens of fs in order to explain the measured I–V characteristics. Specifically, for the results in Fig. 6, $\tau_{\mathrm{LO}}=30$ fs. While such a short phonon lifetime may seem unexpected for bulk GaN, it is reasonable in GaN heterostructures, where LO phonon interactions with interfaces create additional decay channels that enable rapid conversion into heat-conducting modes. Matulionis et al. (2010); Liberis et al. (2014) The impact of $\tau_{\mathrm{LO}}$ on the output current at $V_{\mathrm{GS}}=1$ V and $V_{\mathrm{DS}}=10$ V is shown in Fig. 7. As expected, increasing $\tau_{\mathrm{LO}}$ beyond 30 fs reduces the drain current. Conversely, decreasing $\tau_{\mathrm{LO}}$ from 30 fs to 1 fs yields a $\sim$30% increase in drain current, indicating that even at ultrashort $\tau_{\mathrm{LO}}$ the device performance remains constrained by the hot-LO-phonon bottleneck.
Figure 6: Simulated output curves with 30 fs hot LO phonon lifetime. Symbols are experimental data from Marino et al. Marino et al. (2010)
@@
The role of LO phonon heating is further illustrated in Fig. 8(a), which presents the simulated output characteristics at $V_{\mathrm{GS}}=1~\mathrm{V}$ under four different conditions. The simulation with $\tau_{\mathrm{LO}}=30$ fs is the baseline result (labeled as “All”), which matches well with the measured data shown in Fig. 6. When LO phonon heating with $\tau_{\mathrm{LO}}=30$ fs is considered but acoustic phonon heating is ignored (result labeled as “LO”), the drain current merely increases by 5% at $V_{\mathrm{GS}}=1$ V and $V_{\mathrm{DS}}=10$ V, compared to the baseline case. This observation reinforces that LO phonon heating alone introduces substantial performance degradation even in the absence of acoustic-phonon-induced self-heating. When we assume an instantaneous decay of LO phonons into acoustic modes, effectively removing the hot-LO phonon contribution and considering only longitudinal and transverse acoustic phonon heating (labeled as “LA/TA”), $I_{\mathrm{D}}$ increases by approximately 30% at $V_{\mathrm{GS}}=1$ V and $V_{\mathrm{DS}}=10$ V. In contrast, eliminating both hot LO and hot acoustic phonons in the device (labeled as “None”) would increase the device current by up to 50%, suggesting that phonon-engineering strategies could, in principle, enhance intrinsic device performance. However, we are not aware of any studies demonstrating LO-phonon lifetimes below $30\pm 15$ fs in GaN heterostructures, indicating that such aggressive lifetime reduction has not yet been realized experimentally.